write an equation of the line parallel to 5y = 3x + 12 that has a y-intercept of 2
Answer by TimothyLamb(4379)

slope:
5y = 3x + 12
y = (3/5)x + 12/5
m = 3/5
---
y = (3/5)x + 2
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---